# Motivic Homotopy Groups of Spheres and Free Summands of Stably Free Modules

### v2

@@ 5. Application to maps of Stiefel varieties and stably free modules
For any $r\in\{1,\dots,n\}$, there is a projection map $\rho:V_{r}(\mathbb{A}^{n})\to V_{1}(\mathbb{A}^{n})$. The problem of determining whether $\rho$ has a right inverse (alternatively termed “a section” since $\rho$ is a fibre bundle) was studied in [9]. Using Theorem 4.5, we can settle all cases of this problem over $\bar{\mathbb{Q}}$.
We recall the definition of the James numbers $b_{q}$, which are also known as the Atiyah–Todd numbers $M_{q}$. Definitions are given in [16], [4], and these definitions are proved to be equal in [1]. Explicitly, they are described by their $p$-adic valuations, $v_{p}$, for all primes $p$:
@@
Theorem 5.1.
Let $n$ be a positive integer and suppose $2\leq r\leq n-2$. Then
has a right inverse if and only if $b_{r}\mid n$
@@
If $b_{r}\nmid n$, then there cannot be a right inverse by virtue of [20, Thm. 6.5]. Note that this result disallows right inverses even in the $\mathbb{A}^{1}$-homotopy category $\text{\bf H}(\bar{\mathbb{Q}})$.
Let us therefore suppose that $b_{r}\mid n$, and try to construct a right inverse. The case of $r=2$ follows from a well known construction of Bass, first published to our knowledge in [20, Prop. 2.2(b)] as a construction of a free summand in a projective module.
@@
Theorem 4.5 applies with (n2,n)(n-2,n) playing the part of $(d,e)$ and (r1,n1)(r-1,n-1) playing the part of $(r,n)$. Inequality 2 of Theorem 4.5 required $r<n/2$ in this case. We deduce that the realization map
$\pi_{n-2+n\alpha}(V_{r-1}(\mathbb{A}^{n-1}))\to\pi_{2n-2}(W_{r-1}(\mathbb{C}^{n-1})$
is injective. We now use [9, Prop. 7.3], which tells us that $\rho$ has a right inverse if the corresponding map of complex Stiefel manifolds
@@
has a right inverse. This has a right inverse by [1, Thm. 1.1]. ∎
Corollary 5.2.
$P\oplus R\cong R^{n}$
$P\cong Q\oplus R^{r-1}.$
Proof.
This follows from Theorem 5.1 and [9, Prop. 4.4]. ∎
@@ Appendix A Homological algebra modulo uniquely divisible subgroups
On the $E_{2}$-page and later, there are no nonzero differentials arriving at or emanating from the $(3,0)$-group, which must be 0 by the $E_{\infty}$-page, so that we deduce that the upper sequence in (11) is exact at $\ker(\phi_{3})$. The terms in this sequence, other than $\ker(\phi_{3})$, are known to be uniquely divisible, so that the $5$-lemma implies that $\ker(\phi_{3})$ is also uniquely divisible.

### v3

@@ 5. Application to maps of Stiefel varieties and stably free modules
For any $r\in\{1,\dots,n\}$, there is a projection map $\rho:V_{r}(\mathbb{A}^{n})\to V_{1}(\mathbb{A}^{n})$. The problem of determining whether $\rho$ has a right inverse (alternatively termed “a section” since $\rho$ is a fibre bundle) was studied in [undefi]. Using Theorem 1.6, we can settle all cases of this problem over $\bar{\mathbb{Q}}$.
We recall the definition of the James numbers $b_{q}$, which are also known as the Atiyah–Todd numbers $M_{q}$. Definitions are given in [undefs], [undefc], and these definitions are proved to be equal in [undef]. Explicitly, they are described by their $p$-adic valuations, $v_{p}$, for all primes $p$:
@@
This brings us to our last main theorem, which we restate for the benefit of the reader. See 1.7
@@
If $b_{r}\nmid n$, then there cannot be a right inverse by virtue of [Raynaud1968, Thm. 6.5]. Note that this result disallows right inverses even in the $\mathbb{A}^{1}$-homotopy category $\text{\bf H}(\bar{\mathbb{Q}})$.
Let us therefore suppose that $b_{r}\mid n$, and try to construct a right inverse. The case of $r=2$ follows from a well known construction of Bass, first published to our knowledge in [Raynaud1968, Prop. 2.2(b)] as a construction of a free summand in a projective module.
@@
Theorem 1.6 applies with (n2,n)(n-2,n) playing the part of $(d,e)$ and (r1,n1)(r-1,n-1) playing the part of $(r,n)$. Inequality 2 of Theorem 1.6 requires $r<n/2$ in this case. The conclusion of the theorem is that the realization map
$\pi_{n-2+n\alpha}(V_{r-1}(\mathbb{A}^{n-1}_{\bar{\mathbb{Q}}}))\to\pi_{2n-2}(W_{r-1}(\mathbb{C}^{n-1}))$
is injective. We now use [undefi, Prop. 7.3], which tells us that $\rho$ has a right inverse if the corresponding map of complex Stiefel manifolds
@@
has a right inverse. This has a right inverse by [undef, Thm. 1.1]. ∎
@@ Appendix A Homological algebra modulo uniquely divisible subgroups
On the $E_{2}$-page and later, there are no nonzero differentials arriving at or emanating from the $(3,0)$-group, which must be 0 by the $E_{\infty}$-page, so that we deduce that the upper sequence in (13) is exact at $\ker(\phi_{3})$. The terms in this sequence, other than $\ker(\phi_{3})$, are uniquely divisible by hypothesis, so that the five-lemma implies that $\ker(\phi_{3})$ is also uniquely divisible.